OpenStudy (swissgirl):

okkkk last one hopefully R, where (x,y)R(z,w) iff \( x+z \leq y+w \) on the set R x R

OpenStudy (swissgirl):

its not symmetric obviously

OpenStudy (swissgirl):

Its reflexive

OpenStudy (swissgirl):

and asual transitivity is throwing me off

OpenStudy (swissgirl):

how wld u show transitivity?

OpenStudy (experimentx):

1 + 2 <= 1 + 3 1 + 3 <= 1 + 2 (false)

OpenStudy (experimentx):

1 + 2 <= 1+ 3 <= 1 + 4 1 + 2 <= 1 + 4 (true) 1 + 2 <= 1 + 2 (true)

OpenStudy (swissgirl):

huhhh i didnt follow that

OpenStudy (swissgirl):

i mean i see how its symmetric

OpenStudy (swissgirl):

but not transitive

OpenStudy (experimentx):

symmetric false transitive true reflexive true

OpenStudy (swissgirl):

ya but can u show how its transitive

OpenStudy (swissgirl):

cuz i didnt follow ur example

OpenStudy (experimentx):